# Evaluating the Posterior Sampling Ability of Plug&Play Diffusion Methods in Sparse-View CT

### v1

@@ Abstract
Plug&Play (PnP) diffusion models are state-of-the-art methods in computed tomography (CT) reconstruction. Such methods usually consider applications where the sinogram contains a sufficient amount of information for the posterior distribution to be peaked, and consequently are evaluated using image-to-image metrics such as PSNR/SSIM. Instead, we are interested in reconstructing compressible flow images from sinograms having a small number of projections, which results in a posterior distribution no longer peaked or even multimodal. Thus, in this paper, we aim at evaluating the approximate posterior of PnP diffusion models and introduce two posterior evaluation criteria. We quantitatively evaluate three PnP diffusion methods on three different datasets for several numbers of projections. We surprisingly find that, for each method, the approximate posterior deviates from the true posterior when the number of projections decreases.
@@ 1 Introduction
Diffusion models learn the prior of an underlying data distribution, which allows to generate new samples [1, 2, 3, 4, 5, 6]. Plug&Play (PnP) diffusion models [4, 7, 8, 9, 10, 11, 12, 13] employ such prior-encoding models to solve inverse problems. In this paper, we focus on the Sparse-View Computed Tomography (SVCT) [14] measurement model :
$\mathbf{y}_{p}=\mathbf{H}_{p}\mathbf{x}+\boldsymbol{\epsilon},$ (1)
where $\mathbf{y}_{p}\in\mathbb{R}^{m_{p}}$ is the measured sinogram with $p$ projections, $\mathbf{H}_{p}$ is a discretized Radon transform matrix of size $m_{p}\times n$ corresponding to a parallel beam setting, $\mathbf{x}\in\mathbb{R}^{n}$ is the image and $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\sigma_{y}^{2}\mathbf{I})$ is the measurement noise. At test time, only $\mathbf{y}_{p}$, $\mathbf{H}_{p}$, $\sigma_{y}^{2}$ and the prior-encoding diffusion model are known. In practice, $\mathbf{H}_{p}$ and $\sigma_{y}^{2}$ depend on the tomograph settings. For this reason, we focus on PnP diffusion models that avoid any extra learning stage specific to the measurement model and/or its parameters, as opposed to conditional diffusion models [15, 16, 17].
The vast majority of PnP diffusion models for SVCT [18, 19, 20, 10, 9] consider applications where the measured sinogram contains a sufficient amount of information for the posterior distribution $p(\mathbf{x}|\mathbf{y}_{p})$ to be relatively peaked (see Fig. 1). In this case, a point estimation of $\mathbf{x}$ is sufficient and image-to-image metrics such as PSNR/SSIM [21] may be employed to evaluate the performances of a given approach. Instead, this paper is motivated by the reconstruction of compressible flow images from measured sinograms having a small number of projections, whether for limited optical access or complexity of the setup, as it is the case of digital holographic interferometry [22], where $p=6$. In this context, the posterior distribution $p(\mathbf{x}|\mathbf{y}_{p})$ is no longer peaked and may even be multimodal (see Fig. 1). In this case, a point estimation of $\mathbf{x}$ is no longer sufficient, and being able to sample the true posterior distribution $p(\mathbf{x}|\mathbf{y}_{p})$ is fundamental. PnP diffusion methods are approximate posterior samplers. However, to the best of our knowledge, their ability to sample the true posterior, in SVCT, has not been evaluated as of yet.
@@
(i) We derive two properties that a "good" approximate posterior should satisfy, one in image space and the other in measurement space.
@@ 2 Related work
State-of-the-art SVCT methods [11, 10, 18, 19, 23, 24, 9, 25] are essentially developed and evaluated to deal with cases where the posterior is peaked, which often corresponds to a large number of projections. Let us highlight that the term Ultra Sparse-View CT (USVCT) [26, 19, 18, 27] has emerged to perform CT reconstruction from highly sparse sinograms. However, these works are still limited to sinograms with at least $10$ projections, a regime where the posterior is, in the considered applications, still peaked. To the best of our knowledge, this paper is the first attempt to evaluate PnP diffusion methods where the measured sinogram does not contain sufficient information for the posterior distribution to be peaked (see Fig.1).
All of the above methods report their results in terms of PSNR and/or SSIM [21]. As these metrics focus on image-to-image comparisons, an actual true posterior sample may obtain a low score when the posterior is not peaked. Thus they are not well suited to evaluate the ability of a given method to sample the true posterior. On some other inverse problems, e.g. inpainting [28, 11, 12], super-resolution [15, 12], generative metrics such as Fréchet Inception Distance (FID) [29] and Kernel Inception Distance (KID) [30] are used to evaluate the distribution of reconstructed samples. To the best of our knowledge, such generative metrics have not yet been employed in SVCT and this paper is the first attempt to do so.
@@ 3 Background
Diffusion models [2, 4, 3, 6] aim at sampling from a prior distribution $p_{0}$, by defining a generative process that gradually transforms a sample from a known distribution $p_{T}$ into a sample from $p_{0}$. To do so, the following reverse-time Itô Stochastic Differential Equation (SDE) [4] is solved:
@@
where $\mathbf{f}(\mathbf{x},t)$ is the drift function, $g(t)$ is the diffusion coefficient, $\bar{\mathbf{w}}$ is the standard Wiener process flowing backward in time, and $p_{t}$ is a perturbed version of the prior distribution $p_{0}$: $p_{t}(\mathbf{x}^{\prime})=\int p_{t}(\mathbf{x}^{\prime}|\mathbf{x})p_{0}(\mathbf{x})\text{d}\mathbf{x}$ (with $p_{t}(\mathbf{x}^{\prime}|\mathbf{x})$ being the perturbation kernel). Throughout the paper, we consider the variance-exploding (VE) SDE settings111 VE-SDE settings: $\textbf{f}(\mathbf{x},t)=\mathbf{0}$, $g^{2}(t)=\frac{\text{d}}{\text{d}t}\sigma^{2}(t)$, $p_{t}(\mathbf{x}^{\prime}|\mathbf{x})=\mathcal{N}(\mathbf{x},\sigma^{2}(t)\mathbf{I})$ and $\sigma(t)=\sigma_{\text{min}}(\sigma_{\text{max}}/\sigma_{\text{min}})^{t}$. In practice, the time-dependent prior score function $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$ is approximated by a neural network $\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{x},t)$ trained over a set of $K$ example images $\left\{\mathbf{x}^{(k)}\right\}_{k=1}^{K}$ using Denoising Score-Matching [31]. In the rest of the paper, we will refer to $\mathbf{s}_{\boldsymbol{\theta}}$ as the time-dependent prior score network. Thus, to (approximately) sample from $p_{0}$, the reverse-time SDE Eq. (2) is simulated with $\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{x},t)$ instead of $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$.
PnP diffusion models [11, 10, 12] aim at sampling from the posterior distribution $p_{0}(\mathbf{x}|\mathbf{y})$ while avoiding any extra learning stage specific to the measurement model and/or its parameters by leveraging a prior-encoding denoiser represented by the score network $\mathbf{s}_{\boldsymbol{\theta}}$. To do so, the following reverse-time SDE is employed:
@@
$\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x}|\mathbf{y})$ $=\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})+\nabla_{\mathbf{x}}\log p_{t}(\mathbf{y}|\mathbf{x}),$ (4)
$\approx\,\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)\quad\,\,\,+\nabla_{\mathbf{x}}\log p_{t}(\mathbf{y}|\mathbf{x}).$ (5)
Because of its dependence w.r.t. $t$, $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{y}|\mathbf{x})$ is not tractable. To circumvent this problem, [10, 32] applies projections onto the measurement subset, [9, 19, 33] solve an optimization problem and [11, 12, 10] propose analytical approximations. From this point of view, PnP diffusion methods are "approximately" true posterior samplers. In this paper, we focus on evaluating how good these approximations are, on SVCT, and consider three state-of-the-art methods MCG [10], DPS [11], and $\Pi$G [12]. Their analytical approximations of $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{y}|\mathbf{x})$ are summarized in Tab. 1.
@@
Table 1: Approximations of $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}_{p}|\mathbf{x}_{t})$ for three state-of-the-art PnP diffusion methods on the SVCT measurement model (Eq. (1)). In the table, $\alpha(\mathbf{x}_{t},\mathbf{y}_{p})$ are handcrafted weights (see Sec. 5.3), $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t})=\mathbf{x}_{t}+\sigma_{t}^{2}\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)$ is Tweedie’s denoised prediction and $\mathbf{H}_{p}^{{\dagger}}$ is a pseudo-inverse of $\mathbf{H}_{p}$. In this paper, we focus on evaluating how good these approximations are when the posterior is not peaked or even multimodal.
4 Posterior evaluation criteria
In this section, we introduce two simple criteria that will allow us to compare the approximate posterior, i.e. samples from a PnP diffusion method, to the true unknown posterior.
@@ 4 Posterior evaluation criteria
where Eq. (7) is derived from the result $\mathbb{E}_{\mathbf{y}_{p}\sim p_{\text{Y}_{p}|\text{X}}}\left[\|\mathbf{y}_{p}-\mathbf{H}_{p}\mathbf{x}\|_{2}^{2}\right]=m_{p}\sigma_{y}^{2}$. Let us highlight that Eq. (6) and Eq. (7) are independent of the number of projections $p$. The left hand side of both equations requires evaluating an expectation over $p_{\text{Y}_{p}}$. This can be done empirically by successively sampling from the prior $p_{\text{X}}$ and the likelihood $p_{\text{Y}_{p}|\text{X}}$.
@@
In practice, the true posterior $p_{\text{X}|\text{Y}_{p}}$ is unknown, but a good approximate posterior $\overset{\sim}{p}_{\text{X}|\text{Y}_{p}}$ is expected to exhibit the same properties. In this context, replacing $p_{\text{X}|\text{Y}_{p}}$ with $\overset{\sim}{p}_{\text{X}|\text{Y}_{p}}$ in Eq. (6) states that a set of approximate posterior samples (where the conditioning sinograms are sampled from $p_{\text{Y}_{p}}$, see above) $\{\hat{\mathbf{x}}_{p}^{(i)}\}_{i=1}^{I}$ should not be distinguishable from a set of prior samples $\{\mathbf{x}^{(i)}\}_{i=1}^{I}$. Therefore, any classical metric computing a distance between two sets of samples, such as FID, KID or CMMD [36], should be close to zero and independent of the number of projections $p$.
@@
However, this criterion alone is not sufficient to evaluate the gap between $p_{\text{X}|\text{Y}_{p}}$ and $\overset{\sim}{p}_{\text{X}|\text{Y}_{p}}$, as any method essentially ignoring the conditioning sinogram $\mathbf{y}_{p}$ and sampling from the prior would satisfy it. To circumvent this limitation, we also evaluate the second criterion Eq. (7), that can be computed as follows: $\sum_{i=1}^{I}\|\mathbf{y}_{p,i}-\mathbf{H}_{p}\hat{\mathbf{x}}_{p,i}\|_{2}^{2}/(Im_{p}\sigma_{y}^{2})$ where $m_{p}$ is the dimension of sinograms and the expectation over the posterior is reduced to a single posterior sample. In the remainder of the paper, this quantity is referred as the Normalized average Measurement Consistency (NMC). Let us highlight that the NMC should be close to 1. A value close to 0 indicates that the method only "transforms" the conditioning sinogram into an image without denoising it. A value significantly larger than 1 reveals that the method essentially ignores the conditioning sinogram.
@@
In practice, the gap between the approximate posterior and the true posterior has two origins (besides the SDE discretization): the approximation $\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)\approx\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$, and the approximation of $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}_{p}|\mathbf{x}_{t})$ (see Tab. 1). As this paper focuses on evaluating the latter, in the experiments we get rid of the first approximation as follows: each time a prior sample is required, we obtain it by sampling (unconditionally) using $\mathbf{s}_{\boldsymbol{\theta}}$. Consequently, in the experiments, $\mathbf{s}_{\boldsymbol{\theta}}$ is no longer an approximation but the exact time-dependent prior score function, and only the approximation of $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}_{p}|\mathbf{x}_{t})$ is evaluated.
Another advantage of this procedure is that an arbitrary large number of prior images can be sampled, which is very important for the FID, that needs to be computed on large sets of images to be unbiased [37, 36].

### v2

@@ Abstract
Plug&Play (PnP) diffusion models are state-of-the-art methods in computed tomography (CT) reconstruction. Such methods usually consider applications where the sinogram contains a sufficient amount of information for the posterior distribution to be concentrated around a single mode, and consequently are evaluated using image-to-image metrics such as PSNR/SSIM. Instead, we are interested in reconstructing compressible flow images from sinograms having a small number of projections, which results in a posterior distribution no longer concentrated or even multimodal. Thus, in this paper, we aim at evaluating the approximate posterior of PnP diffusion models and introduce two posterior evaluation properties. We quantitatively evaluate three PnP diffusion methods on three different datasets for several numbers of projections. We surprisingly find that, for each method, the approximate posterior deviates from the true posterior when the number of projections decreases.
@@ 1 Introduction
Diffusion models learn the prior of an underlying data distribution, allowing to generate new samples [1, 2, 3, 4, 5, 6]. Plug&Play (PnP) diffusion models [4, 7, 8, 9, 10, 11, 12, 13] employ such prior-encoding models to solve inverse problems. In this paper, we focus on the Sparse-View Computed Tomography (SVCT) [14] measurement model :
$\mathbf{y}_{p}=\mathbf{H}_{p}\mathbf{x}+\boldsymbol{\epsilon}_{p},$ (1)
where $\mathbf{y}_{p}\in\mathbb{R}^{m_{p}}$ is the measured sinogram with $p$ projections, $\mathbf{H}_{p}$ is a discretized Radon transform matrix of size $m_{p}\times n$ corresponding to a parallel beam setting, $\mathbf{x}\in\mathbb{R}^{n}$ is the image and $\boldsymbol{\epsilon}_{p}\sim\mathcal{N}(\mathbf{0},\sigma_{y}^{2}\mathbf{I}_{m_{p}})$ is the measurement noise.
At test time, only $\mathbf{y}_{p}$, $\mathbf{H}_{p}$, $\sigma_{y}^{2}$ and the prior-encoding diffusion model are known. In practice, $\mathbf{H}_{p}$ and $\sigma_{y}^{2}$ depend on the tomograph settings. For this reason, we focus on PnP diffusion models that avoid any extra learning stage specific to the measurement model and/or its parameters, as opposed to conditional diffusion models [15, 16, 17].
The majority of PnP diffusion models for SVCT [18, 19, 20, 10, 9] consider applications where the measured sinogram contains a sufficient amount of information for the posterior distribution $p(\mathbf{x}|\mathbf{y}_{p})$ to be peaked, i.e. concentrated around a single mode (see Fig. 1). In this case, a point estimation of $\mathbf{x}$ is sufficient to evaluate such posterior and consequently image-to-image metrics such as PSNR/SSIM [21] may be employed to evaluate the performances of a given approach. Instead, this paper is motivated by the reconstruction of compressible jet flows images from a small number of projections [22], whether for limited optical access or complexity of the setup, as it is the case of digital holographic interferometry [22], where $p=6$. In this context, the posterior distribution $p(\mathbf{x}|\mathbf{y}_{p})$ is no longer peaked and may even be multimodal (see Fig. 1). In this case, a point estimation of $\mathbf{x}$ is no longer sufficient to evaluate such posterior, therefore sampling the true posterior distribution $p(\mathbf{x}|\mathbf{y}_{p})$ is fundamental. PnP diffusion methods are approximate posterior samplers. However, to the best of our knowledge, their ability to sample the true posterior, in SVCT, has not been evaluated as of yet.
@@
(i) We derive two properties that a “good" approximate posterior should satisfy, one in image space and the other in measurement space.
@@ 2 Related work
State-of-the-art (SOTA) SVCT methods [11, 10, 18, 19, 23, 24, 9, 25] are essentially developed and evaluated to deal with cases where the posterior is peaked, which often corresponds to a large number of projections. Let us highlight that the term Ultra Sparse-View CT (USVCT) [26, 19, 18, 27] has emerged to perform CT reconstruction from highly sparse sinograms. However, these works are still limited to sinograms with at least $10$ projections, a regime where the posterior is, in the considered applications, still peaked. To the best of our knowledge, this paper is the first attempt to evaluate PnP diffusion methods where the measured sinogram does not contain sufficient information for the posterior distribution to be peaked (see Fig.1).
The above methods report their results in terms of PSNR and/or SSIM [21]. As these metrics focus on image-to-image comparisons, a true posterior sample may obtain a low score when the posterior is not peaked. Thus, they are not well suited to evaluate the ability of a given method to sample the true posterior. In some other inverse problems, e.g. inpainting [28, 11, 12], super-resolution [15, 12], generative metrics such as Fréchet Inception Distance (FID) [29] and Kernel Inception Distance (KID) [30] are used to evaluate the distribution of reconstructed samples. To the best of our knowledge, such generative metrics have not yet been employed in SVCT and this paper is the first attempt to do so.
@@ 3 Background
Diffusion models [2, 4, 3, 6] aim at sampling from a prior distribution $p_{0}$, by defining a generative process that gradually transforms a sample from a known distribution $p_{T}$ into a sample from $p_{0}$. To do so, the following reverse-time Stochastic Differential Equation (SDE) [4] is solved:
@@
where $\mathbf{f}(\mathbf{x},t)$ is the drift function, $g(t)$ is the diffusion coefficient, $\bar{\mathbf{w}}$ is the standard Wiener process flowing backward in time, and $p_{t}$ is a perturbed version of the prior distribution $p_{0}$: $p_{t}(\mathbf{x}^{\prime})=\int p_{t}(\mathbf{x}^{\prime}|\mathbf{x})p_{0}(\mathbf{x})\text{d}\mathbf{x}$, with $p_{t}(\mathbf{x}^{\prime}|\mathbf{x})$ being the perturbation kernel. Throughout the paper, we consider the variance-exploding (VE) SDE settings111 VE-SDE settings: $\textbf{f}(\mathbf{x},t)=\mathbf{0}$, $g^{2}(t)=\frac{\text{d}}{\text{d}t}\sigma^{2}(t)$, $p_{t}(\mathbf{x}^{\prime}|\mathbf{x})=\mathcal{N}(\mathbf{x},\sigma^{2}(t)\mathbf{I})$ and $\sigma(t)=\sigma_{\text{min}}(\sigma_{\text{max}}/\sigma_{\text{min}})^{t}$. In practice, the time-dependent prior score function $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$ is approximated by a neural network $\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{x},t)$ trained over a set of $K$ example images $\left\{\mathbf{x}^{(k)}\right\}_{k=1}^{K}$ using Denoising Score-Matching [31]. In the rest of the paper, we will refer to $\mathbf{s}_{\boldsymbol{\theta}}$ as the time-dependent prior score network. Thus, to (approximately) sample from $p_{0}$, the reverse-time SDE Eq. (2) is simulated with $\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{x},t)$ instead of $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$.
PnP diffusion models [11, 10, 12] aim at sampling from the posterior distribution $p_{0}(\mathbf{x}|\mathbf{y})$ while avoiding any extra learning stage specific to the measurement model and/or its parameters.
To do so, the following reverse-time SDE is employed:
@@
$\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x}|\mathbf{y})$ $=\underbrace{\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})}_{\textstyle\approx\,\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)}+\nabla_{\mathbf{x}}\log p_{t}(\mathbf{y}|\mathbf{x}).$ (4)
Because of its dependence w.r.t. $t$, $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{y}|\mathbf{x})$ is not tractable. To circumvent this problem, [10, 32] applies projections onto the measurement subset, [9, 19, 33] solve an optimization problem and [11, 12, 10] propose analytical approximations. From this point of view, PnP diffusion methods are“approximately" true posterior samplers. In this paper, we focus on evaluating how good these approximations are, on SVCT, and consider three SOTA methods MCG [10], DPS [11], and $\Pi$[12]. Their analytical approximations of $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{y}|\mathbf{x})$ are summarized in Tab. 1.
@@
Table 1: Approximations of $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}_{p}|\mathbf{x}_{t})$ for three SOTA PnP diffusion methods on the SVCT measurement model (Eq. (1)). In the table, $\alpha_{\{\text{DPS},\text{MCG}\}}(\mathbf{x}_{t},\mathbf{y}_{p})$ are handcrafted weights (see Sec. 4.3), $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t})=\mathbf{x}_{t}+\sigma_{t}^{2}\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)$ is Tweedie’s denoised prediction and $\mathbf{H}_{p}^{{\dagger}}$ is a pseudo-inverse of $\mathbf{H}_{p}$. In this paper, we focus on evaluating how good these approximations are when the posterior is not peaked or even multimodal.
4 Posterior evaluation
In this section, we introduce two properties that will allow us to compare the approximate posterior, i.e. samples from a PnP diffusion method, to the true unknown posterior.
@@ 4 Posterior evaluation
where Eq. (6) is derived from the result $\mathbb{E}_{\mathbf{y}_{p}\sim p_{\text{Y}_{p}|\text{X}}}\left[\|\mathbf{y}_{p}-\mathbf{H}_{p}\mathbf{x}\|_{2}^{2}\right]=m_{p}\sigma_{y}^{2}$. Let us highlight that Eq. (5) and Eq. (6) are not dependent on the number of projections $p$. The left hand side of both equations requires evaluating an expectation over $p_{\text{Y}_{p}}$. This can be done empirically by successively sampling from the prior $p_{\text{X}}$ and the likelihood $p_{\text{Y}_{p}|\text{X}}$.
@@
In practice, the true posterior $p_{\text{X}|\text{Y}_{p}}$ is unknown, but a good approximate posterior $\overset{\sim}{p}_{\text{X}|\text{Y}_{p}}$ is expected to exhibit the same properties. In this context, replacing $p_{\text{X}|\text{Y}_{p}}$ with $\overset{\sim}{p}_{\text{X}|\text{Y}_{p}}$ in Eq. (5) states that a set of approximate posterior samples (where the conditioning sinograms are sampled from $p_{\text{Y}_{p}}$, see above) $\{\hat{\mathbf{x}}_{p}^{(i)}\}_{i=1}^{N}$ should not be distinguishable from a set of prior samples $\{\mathbf{x}^{(i)}\}_{i=1}^{N}$. Therefore, any classical metric computing a distance between two sets of samples, such as FID, KID or CMMD [36], should be close to zero and independent of the number of projections $p$. We refer to this distance as Posterior-Prior Similarity (PPS), and $\text{PPS}_{\text{FID}}$ and $\text{PPS}_{\text{CMMD}}$ as the implementation of PPS using FID and CMMD, respectively.
@@
However, this criterion alone is not sufficient to evaluate the gap between $p_{\text{X}|\text{Y}_{p}}$ and $\overset{\sim}{p}_{\text{X}|\text{Y}_{p}}$, as any method essentially ignoring the conditioning sinogram $\mathbf{y}_{p}$ and sampling from the prior would satisfy it. To circumvent this limitation, we also evaluate the second criterion Eq. (6), that can be computed as follows: $\sum_{i=1}^{N}\|\mathbf{y}_{p,i}-\mathbf{H}_{p}\hat{\mathbf{x}}_{p,i}\|_{2}^{2}/({N}m_{p}\sigma_{y}^{2})$ where $m_{p}$ is the dimension of sinograms and the expectation over the posterior is reduced to a single posterior sample. In the remainder of the paper, this quantity is referred as the Normalized average Measurement Consistency (NMC). Let us highlight that the NMC should be close to 1. A value close to 0 indicates that the method only “transforms" the conditioning sinogram into an image without denoising it. A value significantly larger than 1 reveals that the method essentially ignores the conditioning sinogram.
@@
In practice, the gap between the approximate posterior and the true posterior has two origins (besides the SDE discretization): the approximation $\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)\approx\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$, and the approximation of $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}_{p}|\mathbf{x}_{t})$ (see Tab. 1). As this paper focuses on evaluating the latter, in the experiments we get rid of the first approximation as follows: each time a prior sample is required, we obtain it by sampling (unconditionally) using $\mathbf{s}_{\boldsymbol{\theta}}$. Consequently, in the experiments, $\mathbf{s}_{\boldsymbol{\theta}}$ is no longer an approximation but the exact time-dependent prior score function, and only the approximation of $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}_{p}|\mathbf{x}_{t})$ is evaluated. Another advantage of this procedure is that an arbitrary large number of prior images can be sampled, which is very important for the FID, that needs to be computed on large sets of images to be unbiased [37, 36].